OpenStudy (anonymous):

simplify (y^0/2xy^-2*yx^0)^3

OpenStudy (anonymous):

\[\Large \left(\frac{y^0}{2}\cdot xy^{-2}\cdot yx^0\right)^3\] do you mean this ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

then Draw it !!

OpenStudy (anonymous):

\[(y ^{0}/2xy \times yx ^{0})^{3}\] \[(1/2xy ^{-2} \times y)^{3}\] --> y or xto power 0 is equal to 1 \[(y ^{2}/2x \times y)^{3}\]

OpenStudy (anonymous):

\[(y ^{2}/2xy)^{3}\] (y/2x)^3 or \[y ^{3}/8x ^{3}\]